a piece of wire 10 m long, is to be cut into two pieces, one of them two-thirds as long as the other. how should the wire be cut?

Answer by jim_thompson5910(35256)

x = length of first piece
(2/3)x = length of second piece



x + (2/3)x = 10

(3/3)x + (2/3)x = 10

[(3+2)/3]*x = 10

(5/3)x = 10

x = 10*(3/5)

x = 30/5

x = 6

--------------------------------

x = 6
(2/3)*x = (2/3)*6 = 4


The first longer piece is 6 m long. The second shorter piece is 4 m long

Notice how 4/6 = 2/3 and how 6+4 = 10.